# Breaking Abbe’s diffraction limit with harmonic deactivation microscopy

### v2

@@ Discussion
We now turn to discussing signal formation in HADES. The image shows an area of the NbO2 thin film that was thermally damaged before the HADES experiments. In bright-field microscopy, the dark area is visible as well, indicating that material was ablated (Fig. S6a) when damage occurred. The bright spot (dashed profile in Fig. 4) is located at a pointy tip of the thin film to pure substrate transition seen in dark-field microscopy (Fig. S6b). A combination of multiple effects can be the reason for the enhancement THG signal enhancement at this bright spot. First, the third-order susceptibility of the sample can be changed locally, since the NbO2 thin film generates harmonics more efficiently than the pure substrate. Second, near-field confinements of the electric field of the fundamental in subwavelength nanostructures can enhance the HHG efficiency (?). Indeed, the presence of such sub-wavelength nanostructures becomes visible under a scanning electron microscope (Fig. S7). During experimenting, we generally observed an enhancement of the signal when scanning over an edge in the sample, indicating near-field enhancement of the fundamental field, which boosts the efficiency of HHG locally.
@@
The ideas presented can be extended to tighter focusing conditions emitting higher harmonics. The in-depth analysis of the deactivation process measured here and by (?) delivers a three-times resolution improvement of THG microscopy and a 3.8-times resolution improvement for a fifth harmonic generation microscope. Considering a near-infrared excitation at 1800 nm and a high NA objective of 1.35, the expected resolutions for the specific set of wavelengths, polarizations, and material (NbO2) chosen here are 120 nm for THG and 80 nm for the fifth harmonic generation. These estimations are made based on eq. (1) and the experimental findings done in this paper and by (?), which yields a saturation fluence of 1.8 mJ/cm2 for the fifth harmonic. The maximum applicable fluence was herby always low enough to not damage the sample itself, meaning that laser-induced damage ultimately limits this technique. Deepening the understanding of the deactivation process for harmonic generation in materials will lead to an even better resolution by either reducing the saturation fluence, i.e. finding ways to deactivate HHG more efficiently, or by extending the highest applicable peak fluence. This principle is ultimately not limited to the deactivation of HHG by photoexcited carriers, which limits the maximum applicable fluence due to photo-induced damage. Other deactivation methods are also suited for this, i.e. control of the electron-hole recombination with a combination of below band-gap driving wavelengths might also be a suitable pathway (?). Ultimately, resolution in HADES is only limited by the real-space excursion of electrons and holes by the driving laser pulse. This excursion underlies the microscopic currents that radiate HHG, and corresponds to a real-space displacement on the single-digit nanometer scale, the exact value of which depends on electronic structure and laser parameters. As a hypothetical deactivation below this limit would truncate HHG even from the inner region without a deactivation pulse, this excursion length scale sets the ultimate resolution limit for HADES.
Furthermore, other optical nonlinearities such as coherent anti-Stokes Raman scattering (?), stimulated Raman scattering (?), or sum frequency generation (?) can be used to achieve super-resolution with similar concepts. However, the extreme nonlinearity of HHG creates a maximum sensitivity to control pulses. This likely makes HHG and related high-order frequency mixing processes the ideal candidates for resolution improvement by deactivation in particular, and optical emission control and optical switching in general. The wide range of solid-state materials, which have shown harmonic deactivation (?), make HADES an interesting candidate for metrology on dielectric and semiconducting materials, as well as, an important research tool for material science. Purley field effect-driven harmonic deactivation allows even application in biological samples.
@@ Methods
A Ti:Sapphire laser amplifier (Solstice, Spectra-Physics) produces a laser pulse at a central wavelength of 790 nm and with a pulse duration below 72 fs. The repetition rate is set to 2 kHz while the average output power is 7 W. A pulse splitter (BS) sends the laser light into each of the arms. In the excitation arm, the light pumps an optical parametric amplifier (OPA, Topas prime), which emits NIR light centered at 1800 nm. The idler of the parametric process was used to generate the third harmonic since only the idler can reach the desired wavelength. In the modulation arm, the delay of the 800 nm pulse can be adjusted by a mechanical delay stage. This is called the modulating arm. Afterward, a spiral phase plate (PP) converts the intensity profile into a donut shape. Optionally, a second harmonic generation BBO can be introduced after the spiral phase plate to change the central wavelength to 400 nm. Variable attenuators control the average power in each arm independently, while telescopes set the pulse size. A dichroic mirror (DM) combines the two pulses. After this, an objective lens (Obj) focuses this pulse bundle onto the sample (S). The modulation arm uses a combination of a half-wave plate and polarizer for a first attenuation, a second attenuation is achieved by a variable neutral density filter wheel. The excitation arm is attenuated by a series of neutral density filters. A schematic of the experiment can be found in the supplementary information (Fig. S1).
@@
For the deactivation measurements, we removed the spiral phase plate from the modulation arm. A telescope focuses the modulation pulse so that it has an FWHM of 200 µm while the excitation NIR pulse has an FWHM of 40 µm. A tube lens focuses the back-reflected harmonics into a fiber spectrometer (Ocean Insight, QE Pro), that resolves the spectral composition of the harmonics.
@@
A microscope objective (Mitutoyo Plan Apo NIR B) collimates the light from the sample with a magnification of 20x and an NA of 0.4. Subsequently, a tube lens with a focal length of 20 cm images them onto a camera. A pair of Bandpass filters blocked the donut pulse for PSF imaging. They have a central wavelength of 600 nm and a FWHM of 40 nm. The donut pulse was attenuated by an ND filter with an OD of 4.0 to avoid permanent damage. We evaluated the FWHM of the PSF by binarizing the images at a relative intensity of 0.5. Then an ellipse was fitted to the mask created that way. The FWHM represents the average radius of the fitted ellipse.
@@
The images have a pixel size of 1 µm in both directions and are 51 x 51 pixels large. The images were scanned line by line. For the THG microscopy at each pixel 25 laser pulses were measured and each line was repeated 3 times, while for HADES each line was repeated 12 times with 25 laser pulses per pixel to ensure a similar count of photons for each pixel. The microscopy pictures displayed are the averaged values of each measurement, normalized to the maximum and minimum values. For the profile generation, the images were smoothed using a median filter with a pixel radius of 1.5. The profiles were generated with the profile plotter of ImageJ. The profiles were again normalized and interpolated by a spline algorithm for measurements of the FWHM of the features and measurements of the peak separation.
@@ Acknowledgments
Acknowledgments:
We thank Christoph Morscher for help with dark-field images. We thank Antonios Pelekanidis for helping with the step-phase plates. We thank Matthijs Velsink for help with the APD. We thank the AMOLF support department for providing and helping with the measurement software and hardware.

### v3

@@ Discussion
We now turn to discussing signal formation in HADES. The image in Fig. 4 shows an area of the NbO2 thin film that was thermally damaged prior to the HADES experiments. In bright-field microscopy, the dark area is visible as well, indicating that material was ablated (Fig. S4a) when damage occured. The bright spot (dashed profile in Fig. 4) is located at a pointy tip of the thin film to pure substrate transition seen in dark-field microscopy (Fig. S4b). A combination of multiple effects can be the reason for the enhancement THG signal enhancement at htis bright spot. First, the third-order susceptibility of the sample can be changed locally, since the NbO2 thin film generates harmonics more efficiently than the pure substrate. Second, near-field confinements of the electric field of the fundamental in subwavelength nanostructures can enhance the HHG efficiency (?). Indeed, the presence of such sub-wavelength nanostructures becomes visible under a scanning electron microscope (Fig. S5). During experimenting, we generally observed an enhancement of the signal when scanning over an edge in the sample, indicating near-field enhancement of the fundamental field, which boosts the efficiency of HHG locally.
@@
The in-depth analysis of the deactivation process measured here and by (?) delivers a three-time resolution improvement of THG microscopy and a 3.8-time resolution improvement for a fifth harmonic generation microscope. In conjunction with a near-infrared excitation at 1800 nm and a high NA objective of 1.35, the expected resolutions for the specific set of wavelengths, polarizations, and material (NbO2) chosen here are 120 nm for THG and 80 nm for the fifth harmonic generation. Deepening the understanding of the deactivation process for harmonic generation in materials will lead to an even better resolution by either reducing the saturation fluence, i.e. finding ways to deactivate HHG more efficiently, or by extending the highest applicable peak fluence. This principle is ultimately not limited to the deactivation of HHG by photoexcited carriers, which limits the maximum applicable fluence due to photo-induced damage. Other deactivation methods are also suited for this, i.e. control of the electron-hole recombination with a combination of below band-gap driving wavelengths might also be a suitable pathway (?). Ultimately, resolution in HADES is only limited by the real-space excursion of electrons and holes by the driving laser pulse. This excursion underlies the microscopic currents that radiate HHG, and corresponds to a real-space displacement on the single-digit nanometer scale, the exact value of which depends on electronic structure and laser parameters. As a hypothetical deactivation below this limit would actually truncate HHG even from the inner region without a deactivation pulse, this excursion length scale sets the ultimate resolution limit for HADES.
Furthermore, other optical nonlinearities such as coherent anti-Stokes Raman scattering (?), stimulated Raman scattering (?), or sum frequency generation (?) can be used to achieve super-resolution with similar concepts. However, the extreme nonlinearity of HHG creates a maximum sensitivity to control pulses. This likely makes HHG and related high-order frequency mixing processes the ideal candidates for resolution improvement by deactivation in particular, and optical emission control and optical switching in general.
@@ Methods
A Ti:Sapphire laser amplifier (Solstice, Spectra-Physics) produces a laser pulse at a central wavelength of 790 nm and with a pulse duration below 72 fs. The repetition rate is set to 2 kHz while the average output power is 7 W. A pulse splitter (BS) sends the laser light into each of the arms. In the excitation arm, the light pumps an optical parametric amplifier (OPA, Topas prime), which emits NIR light centered at 1800 nm. The idler of the parametric process was used to generate the third harmonic since only the idler can reach the desired wavelength. In the modulation arm, the delay of the 800 nm pulse can be adjusted by a mechanical delay stage. This is called the modulating arm. Afterwards, a vortex phase plate (PP) converts the intensity profile into a donut shape. Optionally, a second harmonic generation BBO can be introduced after the vortex phase plate to change the central wavelength to 400 nm. Variable attenuators control the average power in each arm independently, while telescopes set the pulse size. A dichroic mirror (DM) combines the two pulses. After this, an objective lens (Obj) focuses this pulse bundle onto the sample (S). The modulation arm uses a combination of a half-wave plate and polarizer for a first attenuation, a second attenuation is achieved by a variable neutral density filter wheel. The excitation arm is attenuated by a series of neutral density filters. A schematic of the experiment can be found in the supplementary information (Fig. S1).
@@
For the deactivation measurements, we removed the vortex phase plate from the modulation arm. A telescope focuses the modulation pulse so that it has an FWHM of 200 µm while the excitation NIR pulse has an FWHM of 40 µm. A tube lens focuses the back-reflected harmonics into a fiber spectrometer (Ocean Insight, QE Pro), that resolves the spectral composition of the harmonics.
@@
A microscope objective (Mitutoyo Plan Apo NIR B) collimates the light from the sample with a magnification of 20x and an NA of 0.4. Subsequently, a tube lens with a focal length of 20 cm images them onto a camera. A pair of Bandpass filters blocked the donut pulse for PSF imaging. They have a central wavelength of 600 nm and a FWHM of 40 nm. The donut pulse was attenuated by an ND filter with an OD of 4.0 to avoid permanent damage. The images were taken
We evaluated the FWHM of the PSF by binarizing the images at a relative intensity of 0.5. Then an ellipse was fitted to the mask created that way. The FWHM represents the average radius of the fitted ellipse.
@@
The images have a pixel size of 1 µm in both directions and are 51 x 51 pixels large. The images were scanned line by line. For the THG microscopy at each pixel 25 laser pulses were measured and each line was repeated 3 times, while for HADES each line was repeated 12 times with 25 laser pulses per pixel. The microscopy pictures displayed are the averaged values of each measurement, normalized to the maximum and minimum values. For the profile generation, the images were smoothed using a median filter with a pixel radius of 1.5. The profiles were generated with the profile plotter of ImageJ. The profiles were again normalized and interpolated by a spline algorithm for measurements of the FWHM of the features and measurements of the peak separation.